OpenStudy (anonymous):

Look at the quadratic equation below. x2 - 2 = -2x2 + 5x What is this quadratic equation written in standard form? A. 3x2 - 5x - 2 = 0 B. x2 + 5x - 2 = 0 C. x2 - 5x - 2 = 0 D. 3x2 + 5x - 2 = 0

OpenStudy (aravindg):

TAKE ALL TERMS TO LHS

OpenStudy (aravindg):

@nathaly95 ?

OpenStudy (anonymous):

Yeah and after that?

OpenStudy (aravindg):

add like terms

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Do you get how to do it? Don't be shy we are internet strangers we don't care.

OpenStudy (aravindg):

what did u get?

OpenStudy (aravindg):

........?

OpenStudy (kirbykirby):

x2 - 2 = -2x2 + 5x x2 - 2 + 2x2 - 5x = 0 transferring everything to the left hand side 3x2 - 5x - 2 = 0 adding like-terms

OpenStudy (aravindg):

@kirbykirby we do not appreciate giving away direct answers like that .This is against openstudy COC

OpenStudy (kirbykirby):

Oh sorry. I thought maybe after an hour they must have been very stuck

OpenStudy (aravindg):

But thats not the case here , if she had actually worked on it then she would hav immediately shown her working to us

OpenStudy (aravindg):

@kirbykirby please be careful next time ! happy studying!

OpenStudy (kirbykirby):

All right then. Have a good night (for me anyway!)

OpenStudy (kirbykirby):

Not sure what time zone you're in ;)

OpenStudy (aravindg):

its morning here :)

OpenStudy (kirbykirby):

AH I see then. Well have a good day then

OpenStudy (aravindg):

thanks